# Ehrhart quasi-polynomials via Barnes polynomials and discrete moments of parallelepipeds

### v1

@@ 5 The unit cube in dimension d
We may notice that the case $j=0$ of identity (61) was already covered by Corollary 5, in the special case of unimodular polytopes. The case $j=d$ of identity (61) seems interesting, and using $B_{0}({\mathbf{a}}_{v})=\frac{1}{\prod_{k=1}^{d}\langle w_{k}(v),z\rangle}$, tells us that
@@ 6 Proof of Theorem 1
[9, Theorem 3.4]. We use the notation $\mathcal{K}_{v}-v:=\mathcal{K}_{0}(v)$, the translation of the cone $\mathcal{K}_{v}$ to the origin. We observe that $t\mathcal{K}_{0}(v)=\mathcal{K}_{0}(v)$ for all $t>0$, a trivial but useful fact. Hence, when we dilate $\mathcal{K}_{v}$ by any $t>0$, all of its edge vectors remain the same, and only its apex $v$ gets dilated:
@@
where we replaced $a_{k}$ by $\langle w_{k}(v),z_{0}\rangle$. In other words, we used the definition of the Barnes polynomial, but with the vector ${\mathbf{a}}_{v}:=\Big(\langle w_{1}(v),z_{0}\rangle,\dots,\langle w_{d}(v),z_{0}\rangle\Big)^{T}$. To recap, using (69), we may sum the latter contributions (73), one from each of the vertex tangent cones, to obtain:
@@ 7 Proof of Theorem 2
$\frac{1}{m!}\mu_{m}(t{\mathcal{P}},z_{0})=(-1)^{d}\sum_{v\in V}\sum_{k=0}^{d+m}\sum_{j=0}^{k}\tfrac{1}{k!(d+m-k)!}\binom{k}{j}B_{j}({\mathbf{a}}_{v})t^{k-j}\langle v,z_{0}\rangle^{k-j}\sum_{q\in\mathbb{Z}^{d}-tv\;(\hskip-2.84544pt\bmod\Pi_{v})}{\langle q,z_{0}\rangle}^{d+m-k},$ (88)
@@
$(-1)^{d}\sum_{v\in V}\sum_{j=0}^{d+m-r}\frac{1}{(d+m-r-j)!j!r!}B_{j}({\mathbf{a}}_{v})\langle v,z_{0}\rangle^{r}\sum_{q\in\mathbb{Z}^{d}-tv\;(\hskip-2.84544pt\bmod\Pi_{v})}{\langle q,z_{0}\rangle}^{d+m-r-j}$ (89)
$=\frac{(-1)^{d}}{(d+m)!}\sum_{v\in V}\sum_{j=0}^{d+m-r}\binom{d+m}{r,j}B_{j}({\mathbf{a}}_{v})\langle v,z_{0}\rangle^{r}\sum_{q\in\mathbb{Z}^{d}-tv\;(\hskip-2.84544pt\bmod\Pi_{v})}{\langle q,z_{0}\rangle}^{d+m-r-j},$ (90)
valid for almost all $z_{0}\in\mathbb{C}^{d}$.
To prove part (c), we simply differentiate the latter expression in $t$:
$\frac{d}{dt}d_{r+1}(t)=\frac{d}{dt}\left(\tfrac{(-1)^{d}m!}{(d+m)!}\sum_{v\in V}\sum_{j=0}^{d+m-r}\binom{d+m}{r,j}B_{j}({\mathbf{a}}_{v})\langle v,z\rangle^{r}\sum_{q\in\mathbb{Z}^{d}-tv\;(\hskip-2.84544pt\bmod\Pi_{v})}{\langle q,z\rangle}^{d+m-r-j}\right)$
$=(-1)^{d}m!\sum_{v\in V}\langle v,z\rangle^{r+1}\sum_{j=0}^{d-r-1}\tfrac{(d+m-r-j)}{r!j!(d+m-r-j)!}B_{j}({\mathbf{a}}_{v})\sum_{q\in(\Pi_{v}+tv)\cap\mathbb{Z}^{d}}{\Big(\langle q,z\rangle-t\langle v,z\rangle\Big)}^{d+m-r-j-1}$
@@
8 Proof of Theorem 3: vanishing identities for integer simple polytopes
@@
$\sigma_{t{\mathcal{P}}}(xz_{0})$ $=(-1)^{d}\sum_{v\in\rm{V}}\sigma_{\Pi_{v}}(xz_{0})\sum_{k\geq 0}B_{k}(t\langle v,z_{0}\rangle,{\mathbf{a}}_{v})\frac{x^{k-d}}{k!}.$
To prove part (a), we begin by noting that the left-hand-side of the latter identity is an entire function of $x$. We therefore know that for each fixed $n<0$, the coefficient of any $x^{n}$ on the right-hand-side must be identically zero. We next expand the exponential function $e^{x\langle z_{0},p\rangle}$ as a Taylor series as well, obtaining:
$(-1)^{d}\sigma_{t{\mathcal{P}}}(xz_{0})=$ $\sum_{v\in V}\sum_{p\in\Pi_{v}\cap\mathbb{Z}^{d}}e^{x\langle z_{0},p\rangle}\sum_{k\geq 0}B_{k}(t\langle v,z_{0}\rangle,{\mathbf{a}}_{v})\frac{x^{k-d}}{k!}$ (91)
$=\sum_{v\in V}\sum_{p\in\Pi_{v}\cap\mathbb{Z}^{d}}\sum_{n\geq 0}{\langle z_{0},p\rangle}^{n}\frac{x^{n}}{n!}\sum_{k\geq 0}B_{k}(t\langle v,z_{0}\rangle,{\mathbf{a}}_{v})\frac{x^{k-d}}{k!}$ (92)
$=\sum_{v\in V}\sum_{n,k\geq 0}\sum_{p\in\Pi_{v}\cap\mathbb{Z}^{d}}{\langle z_{0},p\rangle}^{n}B_{k}(t\langle v,z_{0}\rangle,{\mathbf{a}}_{v})\frac{x^{k+n-d}}{k!n!}$ (93)
The left-hand side of (91) is an entire function of $x$. Therefore we cannot have any singularities in $x$, so that comparing the coefficient of $x^{m-d}$ on both sides, for each fixed value of $m\in\{0,\dots,d-1\}$, we must get vanishing identities. Precisely, we get the following $d$ identities for each integer polytope in $\mathbb{R}^{d}$, by summing over all $k,n\geq 0$ in (93) with $k+n=m$ fixed:
$0$ $=\sum_{k=0}^{m}\binom{m}{k}\sum_{v\in V}B_{k}\Big(t\langle v,z_{0}\rangle,{\mathbf{a}}_{v}\Big)\sum_{p\in\Pi_{v}\cap\mathbb{Z}^{d}}{\langle p,z_{0}\rangle}^{m-k}.$ (94)
@@
$0$ $=\sum_{k=0}^{m}\binom{m}{k}\sum_{v\in\rm{V}}\sum_{j=0}^{k}\binom{k}{j}B_{j}({\mathbf{a}}_{v})\langle v,z\rangle^{k-j}t^{k-j}\sum_{p\in\Pi_{v}\cap\mathbb{Z}^{d}}{\langle p,z\rangle}^{m-k}$ (95)
$=\sum_{v\in\rm{V}}\sum_{k=0}^{m}\sum_{j=0}^{k}\binom{m}{k}\binom{k}{j}B_{j}({\mathbf{a}}_{v})\langle v,z\rangle^{k-j}t^{k-j}\sum_{p\in\Pi_{v}\cap\mathbb{Z}^{d}}{\langle p,z\rangle}^{m-k}$ (96)
@@
$0$ $=\sum_{v\in\rm{V}}\langle v,z\rangle^{r}\sum_{j=0}^{m-r}\frac{B_{j}({\mathbf{a}}_{v})}{j!}\sum_{p\in\Pi_{v}\cap\mathbb{Z}^{d}}\frac{{\langle p,z\rangle}^{m-r-j}}{(m-r-j)!},$ (97)
@@
9 Proof of Corollary 2: Ehrhart for unimodular polytopes
@@
(of Corollary 2)The unimodularity assumption means that for all vertices $v\in{\mathcal{P}}$, there is exactly one integer point in the fundamental parallelepiped $\Pi_{v}$, namely the origin. Therefore the discrete moments, for each $0\leq k\leq d-1$, degenerate to:
@@ 10 Examples, and some identities for special cases of the Theorems
In dimension $d=2$, Corollary 5 had an immediate consequence, namely (31), which tells us that for any unimodular integer polygon $P\subset\mathbb{R}^{2}$ we have:
@@
In dimension $d=3$, the same consequence (31) of Corollary 5 tells us that for any unimodular integer polytope $P\subset\mathbb{R}^{3}$, we have:
$1=\frac{(-1)^{3}}{3!}\sum_{v\in\rm{V}}B_{3}({\mathbf{a}}_{v}),$
and unpacking $B_{3}({\mathbf{a}}_{v})$ by using (53), we arrive at the identity
$24=\sum_{v\in\rm{V}}\left(\tfrac{\langle w_{1}(v),z\rangle}{\langle w_{2}(v),z\rangle}+\tfrac{\langle w_{1}(v),z\rangle}{\langle w_{3}(v),z\rangle}+\tfrac{\langle w_{2}(v),z\rangle}{\langle w_{1}(v),z\rangle}+\tfrac{\langle w_{2}(v),z\rangle}{\langle w_{3}(v),z\rangle}+\tfrac{\langle w_{3}(v),z\rangle}{\langle w_{1}(v),z\rangle}+\tfrac{\langle w_{3}(v),z\rangle}{\langle w_{2}(v),z\rangle}+\tfrac{1}{2}\right),$
@@ 10.1 Special cases of Theorem 3
We expand on a few of the vanishing identities from Theorem 3. We first note that the particular cases $r=m$ recovers known identities, by Brion and Vergne [12] and that the other identities appear to be new. Precisely, for $r=m$ with $0\leq m\leq d-1$, these $d$ known identities are:
@@
valid for almost all $z\in\mathbb{C}^{d}$. Here we used $B_{0}({\mathbf{a}})=\frac{1}{a_{1}\cdots a_{d}}$, and $\sum_{p\in\Pi_{v}\cap\mathbb{Z}^{d}}1=|\Pi_{v}\cap\mathbb{Z}^{d}|=\operatorname{vol}\Pi_{v}$. The latter identity is a standard fact, which may be arrived using Fourier transforms of polytopes [12] [37, Theorem 5.11].
Regarding the new identities, we first discuss the case $m=1,r=0$, as it presents a special case that is of independent interest: equation (106) below.
@@
Case (m,r)=(m,m1)(m,r)=(m,m-1), with $2\leq m\leq d$. Here Theorem 3 gives us
@@
for all simple integer polytopes, for almost all $z\in\mathbb{C}^{d}$, and for each $1\leq m\leq d$.
Next, we make a simple observation that holds for the family of unimodular polytopes. If $\mathcal{P}$} is a unimodular polytope, then the latter identity (109) enjoys a simplification. For a unimodular polytope, we have $\sum_{p\in\Pi_{v}\cap\mathbb{Z}^{d}}\langle p,z\rangle=0$, because the origin is the only integer point in $\Pi_{v}$. Therefore
@@
for each fixed $1\leq m\leq d$, and for almost all $z\in\mathbb{C}^{d}$.
@@
for almost all $z\in\mathbb{C}^{2}$. We’ve recovered here a known volume identity of Brion [11, Corollary 2] and Lawrence [27] (see also [37, Theorem 7.23] for another proof).
@@ 11 Further remarks and research directions
Conjecture 1.
Fix the dimension $d$, and a half-open integer paralleleppiped $\Pi\subset\mathbb{R}^{d}$. Then for each fixed integer $k$, and fixed $z\in\mathbb{C}^{d}$, there is a polynomial-time algorithm, in the bit-input of the vertices of an integer parallelepiped $\Pi$, for computing the discrete moment $\mu_{k}(\Pi)$.
@@
We may give a trivial bound for the discrete moments $\mu_{n}(\Pi_{v})$ in Theorem 1, and hence for the coefficients of Ehrhart quasi-polynomials by considering $M_{v}:=\max_{p\in\Pi_{v}}\|p\|$. We have
@@
Let $d\geq 2$. We know from Section 2 that $B_{k}(t\langle v,z\rangle,{\mathbf{a}}_{v})\prod_{j=1}^{d}\langle w_{j},z\rangle$ is a nonconstant polynomial in $z_{1},\dots,z_{d}$, for all $k\geq 1$. Picking a generic $z\in\mathbb{C}^{d}$, in the sense of (7), already implies that $z\in\mathbb{C}^{d}$ satisfies $\prod_{j=1}^{d}\langle w_{j},z\rangle\not=0$. Now pick a generic $z\in\mathbb{C}^{d}$ that also belong to the variety defined by
$\bigcap_{k=1}^{d-1}\left\{z\in\mathbb{C}^{d}\ \mid\ B_{k}(t\langle v,z\rangle,{\mathbf{a}}_{v}).\prod_{j=1}^{d}\langle w_{j},z\rangle=0\right\}.$ (112)
This variety gives us extra flexibility to choose points $z\in\mathbb{C}^{d}$ that allow us to omit the discrete moment sums $\mu_{k}(\Pi_{v})$ (for $k=1,2,\dots,d-1$) from our calculations. Due to the flexibility in picking any generic $z\in\mathbb{C}^{d}$ in the theorems, we arrive at the following conclusion:

### v2

@@ 2.3 A d-dimensional cube
We may notice that the case $j=0$ of identity (48) was already covered by Corollary 2, in the special case of smooth polytopes. The case $j=d$ of identity (48) seems interesting, and using $B_{0}({\mathbf{a}}_{v})=\frac{1}{\prod_{k=1}^{d}\langle w_{k}(v),z\rangle}$, tells us that
@@ 3 Proof of Theorem 1
[10, Theorem 3.4]. We use the notation $\mathcal{K}_{v}-v:=\mathcal{K}_{0}(v)$, the translation of the cone $\mathcal{K}_{v}$ to the origin. We observe that $t\mathcal{K}_{0}(v)=\mathcal{K}_{0}(v)$ for all $t>0$, a trivial but useful fact. Hence, when we dilate $\mathcal{K}_{v}$ by any $t>0$, all of its edge vectors remain the same, and only its apex $v$ gets dilated:
@@
where we replaced $a_{k}$ by $\langle w_{k}(v),z_{0}\rangle$. In other words, we used the definition of the Barnes polynomial, but with the vector ${\mathbf{a}}_{v}:=\Big(\langle w_{1}(v),z_{0}\rangle,\dots,\langle w_{d}(v),z_{0}\rangle\Big)^{T}$. To summarize, by using (56) we may sum the latter contributions (60), one from each of the vertex tangent cones, to obtain:
@@ 4 Proof of Theorem 2
$\tfrac{(-1)^{d}}{m!}\mu_{m}(t{\mathcal{P}},z_{0})=\sum_{v\in V}\sum_{k=0}^{d+m}\sum_{j=0}^{k}\tfrac{1}{k!(d+m-k)!}\binom{k}{j}B_{j}({\mathbf{a}}_{v})t^{k-j}\langle v,z_{0}\rangle^{k-j}\sum_{q\in\mathbb{Z}^{d}-tv\;(\hskip-2.84544pt\bmod\Pi_{v})}{\langle q,z_{0}\rangle}^{d+m-k},$ (75)
@@
$\sum_{v\in V}\sum_{j=0}^{d+m-r}\frac{1}{(d+m-r-j)!j!r!}B_{j}({\mathbf{a}}_{v})\langle v,z_{0}\rangle^{r}\sum_{q\in\mathbb{Z}^{d}-tv\;(\hskip-2.84544pt\bmod\Pi_{v})}{\langle q,z_{0}\rangle}^{d+m-r-j}$ (76)
$=\frac{1}{(d+m)!}\sum_{v\in V}\sum_{j=0}^{d+m-r}\binom{d+m}{r,j}B_{j}({\mathbf{a}}_{v})\langle v,z_{0}\rangle^{r}\sum_{q\in\mathbb{Z}^{d}-tv\;(\hskip-2.84544pt\bmod\Pi_{v})}{\langle q,z_{0}\rangle}^{d+m-r-j},$ (77)
valid for almost all $z_{0}\in\mathbb{C}^{d}$. To prove part (c), we simply differentiate the latter expression in $t$:
$\frac{d}{dt}d_{r}(t)=\frac{d}{dt}\left(\tfrac{m!}{(d+m)!}\sum_{v\in V}\sum_{j=0}^{d+m-r}\binom{d+m}{r,j}B_{j}({\mathbf{a}}_{v})\langle v,z\rangle^{r}\sum_{q\in\mathbb{Z}^{d}-tv\;(\hskip-2.84544pt\bmod\Pi_{v})}{\langle q,z\rangle}^{d+m-r-j}\right)$
$=m!\sum_{v\in V}\langle v,z\rangle^{r+1}\sum_{j=0}^{d-r-1}\tfrac{(d+m-r-j)}{r!j!(d+m-r-j)!}B_{j}({\mathbf{a}}_{v})\sum_{q\in(\Pi_{v}+tv)\cap\mathbb{Z}^{d}}{\Big(\langle q,z\rangle-t\langle v,z\rangle\Big)}^{d+m-r-j-1}$
@@
5 Proof of Theorem 3: vanishing identities for rational simple polytopes
@@
$\sigma_{t{\mathcal{P}}}(xz_{0})$ $=(-1)^{d}\sum_{v\in V}\sum_{q\in\mathbb{Z}^{d}-tv\;(\hskip-2.84544pt\bmod\Pi_{v})}\sum_{n,k\geq 0}{\langle q,z_{0}\rangle}^{n}B_{k}(t\langle v,z_{0}\rangle,{\mathbf{a}}_{v})\frac{x^{k+n-d}}{k!n!}$ (78)
To prove part (a), we begin by noting that the left-hand-side of the latter identity is an entire function of $x$. We therefore know that for each fixed $n<0$, the coefficient of any $x^{n}$ on the right-hand-side must be identically zero. We now compare the coefficient of $x^{m-d}$ on both sides, and for each fixed value of $m\in\{0,\dots,d-1\}$, we must get vanishing identities. Precisely, we get the following $d$ identities for each integer polytope in $\mathbb{R}^{d}$, by summing over all $k,n\geq 0$ in (78) with $k+n=m$ fixed:
$0$ $=\sum_{k=0}^{m}\binom{m}{k}\sum_{v\in V}B_{k}\Big(t\langle v,z_{0}\rangle,{\mathbf{a}}_{v}\Big)\sum_{q\in\mathbb{Z}^{d}-tv\;(\hskip-2.84544pt\bmod\Pi_{v})}{\langle q,z_{0}\rangle}^{m-k}.$ (79)
@@
$0$ $=\sum_{k=0}^{m}\binom{m}{k}\sum_{v\in\rm{V}}\sum_{j=0}^{k}\binom{k}{j}B_{j}({\mathbf{a}}_{v})\langle v,z_{0}\rangle^{k-j}t^{k-j}\sum_{q\in\mathbb{Z}^{d}-tv\;(\hskip-2.84544pt\bmod\Pi_{v})}{\langle q,z_{0}\rangle}^{m-k}$ (80)
$=\sum_{v\in\rm{V}}\sum_{k=0}^{m}\sum_{j=0}^{k}\binom{m}{k}\binom{k}{j}B_{j}({\mathbf{a}}_{v})\langle v,z_{0}\rangle^{k-j}t^{k-j}\sum_{q\in\mathbb{Z}^{d}-tv\;(\hskip-2.84544pt\bmod\Pi_{v})}{\langle q,z_{0}\rangle}^{m-k}$ (81)
@@
$0$ $=\sum_{v\in\rm{V}}\langle v,z\rangle^{r}\sum_{j=0}^{m-r}\frac{B_{j}({\mathbf{a}}_{v})}{j!(m-r-j)!}\sum_{q\in\mathbb{Z}^{d}-tv\;(\hskip-2.84544pt\bmod\Pi_{v})}{\langle q,z_{0}\rangle}^{m-r-j}$ (82)
@@
6 Proof of Corollary 5: Ehrhart for smooth polytopes
@@
(of Corollary 5)The smooth polytope assumption means that for all vertices $v\in{\mathcal{P}}$, there is exactly one integer point in the fundamental parallelepiped $\Pi_{v}$, namely the origin. Therefore the discrete moments, for each $0\leq k\leq d-1$, degenerate to:
@@ 7 Examples, and some identities for special cases of the Theorems
In dimension $d=2$, Corollary 8 tells us that for any smooth integer polygon $P\subset\mathbb{R}^{2}$ we have:
@@
In dimension $d=3$, Corollary 8 tells us that for any smooth integer polytope $P\subset\mathbb{R}^{3}$, we have:
$1=\frac{(-1)^{3}}{3!}\sum_{v\in\rm{V}}B_{3}({\mathbf{a}}_{v})=-\frac{1}{6}\sum_{v\in\rm{V}}\frac{1}{8a_{1}(v)a_{2}(v)a_{3}(v)}(s_{1}s_{2}-s_{1}^{3}),$
by using (113) of Appendix A. Using the definitions of the power sums $s_{j}$ and defining $a_{j}(v):=\langle w_{j}(v),z\rangle$, we arrive at the identity
$24$ $=\frac{1}{2}\sum_{v\in\rm{V}}\frac{s_{1}^{3}-s_{1}s_{2}}{a_{1}(v)a_{2}(v)a_{3}(v)}$ (86)
$=\frac{1}{2}\sum_{v\in\rm{V}}\tfrac{\left(\langle w_{1}(v),z\rangle+\langle w_{2}(v),z\rangle+\langle w_{3}(v),z\rangle\right)^{3}-\left(\langle w_{1}(v),z\rangle+\langle w_{2}(v),z\rangle+\langle w_{3}(v),z\rangle\right)\left(\langle w_{1}(v),z\rangle^{2}+\langle w_{2}(v),z\rangle^{2}+\langle w_{3}(v),z\rangle^{2}\right)}{\langle w_{1}(v),z\rangle\langle w_{2}(v),z\rangle\langle w_{3}(v),z\rangle}$ (87)
$=\sum_{v\in\rm{V}}\left(\tfrac{\langle w_{1}(v),z\rangle}{\langle w_{2}(v),z\rangle}+\tfrac{\langle w_{2}(v),z\rangle}{\langle w_{1}(v),z\rangle}+\tfrac{\langle w_{2}(v),z\rangle}{\langle w_{3}(v),z\rangle}+\tfrac{\langle w_{3}(v),z\rangle}{\langle w_{2}(v),z\rangle}+\tfrac{\langle w_{3}(v),z\rangle}{\langle w_{1}(v),z\rangle}+\tfrac{\langle w_{1}(v),z\rangle}{\langle w_{3}(v),z\rangle}+3\right),$ (88)
@@ 7.1 Special cases of Theorem 3
Restricting attention here to integer polytopes, we expand on a few of the vanishing identities from Theorem 3. We first note that the particular cases $r=m$ recovers known identities, by Brion and Vergne [13] and that the other identities appear to be new. Precisely, for $r=m$ with $0\leq m\leq d-1$, these $d$ known identities are:
@@
valid for almost all $z\in\mathbb{C}^{d}$. Here we used $B_{0}({\mathbf{a}})=\frac{1}{a_{1}\cdots a_{d}}$, and $\sum_{p\in\Pi_{v}\cap\mathbb{Z}^{d}}1=|\Pi_{v}\cap\mathbb{Z}^{d}|=\operatorname{vol}\Pi_{v}$. The latter identity is a standard fact from the geometry of numbers [42, Theorem 5.11].
Throughout this section, we fix the dimension $d\geq 2$. Regarding the new identities, we first discuss the case $m=1,r=0$, as it presents a special case that is of independent interest: equation (94) below.
@@
Case (m,r)=(m,m1)(m,r)=(m,m-1), with $1\leq m\leq d-1$. Here Theorem 3, for the special case of integer simple polytopes, gives us
@@
for all simple integer polytopes, for almost all $z\in\mathbb{C}^{d}$.
Next, we make a simple observation that holds for the family of smooth polytopes. If $\mathcal{P}$} is a smooth polytope, then the latter identity (97) enjoys a simplification. For a smooth polytope, we have $\sum_{p\in\Pi_{v}\cap\mathbb{Z}^{d}}\langle p,z\rangle=0$, because the origin is the only integer point in $\Pi_{v}$. Therefore
@@
for each fixed $1\leq m\leq d-1$, and for almost all $z\in\mathbb{C}^{d}$.
@@
for almost all $z\in\mathbb{C}^{2}$. We’ve recovered here a known volume identity of Brion [12, Corollary 2] and Lawrence [31] (see also [42, Theorem 7.23] for a Fourier-style proof).
@@ 8 Further remarks and research directions
Question 1.
Fix the dimension $d$, and fix a half-open integer paralleleppiped $\Pi\subset\mathbb{R}^{d}$, defined by its edges $w_{1},\dots,w_{d}$. Fix also a positive integer $k$ and $z\in\mathbb{C}^{d}$. Let $N$ be the bit input of the edges $w_{1},\dots,w_{d}$. Is there an $N^{O(d)}$-time algorithm for computing the discrete moment $\mu_{k}(\Pi)$?
Barvinok’s algorithm [6] guarantees a complexity bound of $N^{O(d\log d)}$ for general rational polytopes, where $N$ is the bit input of $\mathcal{P}$}. So in particular Barvinok’s algorithm guarantees a polynomial-time algorithm for computing $\mu_{k}(\Pi)$ in fixed dimension. But with half-open integer parallelepipeds we have a lot more structure, suggesting a slightly better bound.
@@
We may give a trivial bound for the discrete moments $\mu_{n}(\Pi_{v})$ in Theorem 1, and hence for the coefficients of Ehrhart quasi-polynomials by considering $M_{v}:=\displaystyle\max_{p\in\Pi_{v}\cap\mathbb{Z}^{d}}\|p\|$. We have
@@
Let $d\geq 2$. We know from §A that $\left(\prod_{j=1}^{d}\langle w_{j},z\rangle\right)B_{k}\left((t\langle v,z\rangle,{\mathbf{a}}_{v}\right)$ is a nonconstant polynomial in $z_{1},\dots,z_{d}$, for all $k\geq 1$. Picking a generic $z\in\mathbb{C}^{d}$, in the sense of (7), already implies that $z\in\mathbb{C}^{d}$ satisfies $\prod_{j=1}^{d}\langle w_{j},z\rangle\not=0$. Now pick a generic $z\in\mathbb{C}^{d}$ that also belong to the variety defined by
$X:=\bigcap_{k=1}^{d-1}\left\{z\in\mathbb{C}^{d}\ \mid\ \left(\prod_{j=1}^{d}\langle w_{j},z\rangle\right)B_{k}\left((t\langle v,z\rangle,{\mathbf{a}}_{v}\right)=0\right\}.$ (100)
The variety $X$ in (100), as a vanishing set of polynomials in $z\in\mathbb{C}^{d}$, gives us extra flexibility to choose points $z\in\mathbb{C}^{d}$ that allow us to omit the discrete moment sums $\mu_{k}(\Pi_{v})$ (for $k=1,2,\dots,d-1$) from our calculations. Due to the flexibility in picking any generic $z\in\mathbb{C}^{d}$ in the theorems, we arrive at the following conclusion, if we know that the variety (100) is nonempty: